OpenStudy (anonymous):

Multiply or divide as indicated. \[\frac{ 2x }{ 4x+2 } \times \frac{ 14x+7 }{ 6 }\] I'm not sure how to start this problem off

OpenStudy (anonymous):

@JungHyunRan if you're not busy, would you mind helping me with this? Thank you.

OpenStudy (anonymous):

\(\Large \frac{ 2x }{ 4x+2 } \times \frac{ 14x+7 }{ 6 }=\frac{2x(14x+7)}{6(4x+2)}=\frac{x(14x+7)}{3(4x+2)}=\frac{7(2x+1)}{6(2x+1)}=\frac{7}{6}\)

OpenStudy (anonymous):

What level of math are you taking? Or the highest that you've taken because you're really good at this oml and thank you very much

OpenStudy (anonymous):

\(\huge \color{green}{UR\ Welcome}\)